Can someone please properly answer all the questions listed below? I know it's a lot of questions, but it would really mean a lot if someone could help. I'd greatly appreciate it. P.S (I know as per chegg guidlines you can only answer a set amount of questions. But there’s been multiple times where I post like 10-20 question or so and people completely answer them. And others who only answer 1 single question out of everything else. I posted this just before and someone only answered number 1... Please don’t be one of those people.

And also, please (TYPE) everything out.

2- Suppose that the consumption function is C = 100 + 0.5 Y and the planned investment is 50. There is no government in the economy and thus Y = C + S and AE = C + I.

Find the level of total output (or total income) in the equilibrium.

3- Suppose that the government is running a balanced budget and the value of purchases made by the government is 200. The consumption function is C = 200 + 0.6 Yd and planned investment is 100.

a) If the level of aggregate output is 1000, calculate:

- Disposal income

- Aggregate consumption

- Aggregate Saving

- Planned aggregate expenditure

- Unplanned inventory change

b) If the level of aggregate output is 1250, calculate:

- Disposal income

- Aggregate consumption

- Aggregate Saving

- Planned aggregate expenditure

- Unplanned inventory change

Y = C + S + T

Yd = Y – T

Yd = C + S

AE = C + I + G

Unplanned change in inventory: Y – (C + I + G)

4- Knowing that Y = C + S + T and AE = C + I + G, show that in equilibrium the sum of savings and taxes is equal to the sum of planned investment and government purchases.

5- Suppose a consumption function of an economy is C = 500 + 0.8 Yd ,

a) Calculate the government spending multiplier.

b) Calculate the tax multiplier.

c) Calculate the balanced-budget multiplier.

d) If the government increases it purchases for 200, how much the equilibrium outcome will change?

e) If the government decreases taxes for 200, how much the equilibrium outcome will change?

f) If the government increases it purchases for 200 and keep the budget balanced, how much the equilibrium outcome will change?

6- Why is M2 sometimes a more stable measure of money than M1? Explain in your own words using the definitions of M1 and M2. (This question is question 1.3 Chapter 10, Twelfth edition).

7- Name three tools applied by the Federal Reserve to control the interest rate. Briefly explain the mechanism of each one.

8- Explain why the tools you mentioned in question 7 were not effective enough during post 2008 era. How did the Fed manage to raise the interest rate then?

9- What is the difference between fiat money and commodity money?

10- What is the most important asset of banks?

11- What is the main liability of banks?

12- Suppose that the required reserve ratio for a bank is 20 percent. If you deposit $ 1000 in this bank, how much of your money will be kept as cash or reserves at the Federal Reserve?

13- What are the excess reserves of a bank? What do the banks do with the excess reserve?

14- Which one is correct in general? Explain why.

a) The quantity of money demanded (Md) depends postively on the interest rate.

b) The quantity of money demanded (Md) depends negatively on the interest rate.

15- Choose the correct option.

The face value of a 10-year bond issued by the Treasury is the amount of money

a) Paid to the buyer of the bond per year for ten years.

b) Paid to the buyer the buyer after ten years.
Answer #1

2. Consumption C= 100+0.5Y and planned investment I = 50

Then, AE=Y=C+I

or, Y=100+0.5Y+50

or, 0.5Y=150

or, Y=300

3. Y=C+I+G

or, 1000=200+0.6Yd+100+200

or, 0.6Yd=500

or, Yd=833.333

Here Yd is the disposable income.

Aggregate consumption C=200+0.6*833.33 = 200+500 =700

Aggregate savings S=Yd-C = 833.333-700 = 133.333

Planned aggregate expenditure AE=C+I+G = 700+100+200 = 1000

Unplanned inventory change =Y-(C+I+G)=0